OpenStudy (anonymous):

Naomi wants to take yoga classes. She can take classes at the high school for a one-time fee of $20 plus $10 per class. She can join the local gym for $50 and take classes for $5 per class. Which statement is true about the average cost for the two options?

OpenStudy (anonymous):

A)The high school option costs less per class for six classes. B)The gym option costs less per class for six classes. C)The high school option costs less per class for seven classes. D) The gym option costs less per class for seven classes.

OpenStudy (anonymous):

fixing number of classes to a particular amount lets say 6 , then in her high school she has to pay 20a plus 10 per class and we have 6 class, so total amount, 20 + 6*10 = 20 +60 = 80 dollars now in local gym the figure will be 50 + 6*5 = 50 + 30 = 80 dollars so for six classes, both would cost the same , for seven maybe local gym is the better option :D :P